# The Fundamental Limits of Least-Privilege Learning

### v1

@@ Abstract
The promise of least-privilege learning – to find feature representations that are useful for a learning task but prevent inference of any sensitive information unrelated to this task – is highly appealing. However, so far this concept has only been stated informally. It thus remains an open question whether and how we can achieve this goal. In this work, we provide the first formalisation of the least-privilege principle for machine learning and characterise its feasibility. We prove that there is a fundamental trade-off between a representation’s utility for a given task and its leakage beyond the intended task: it is not possible to learn representations that have high utility for the intended task but, at the same time prevent inference of any attribute other than the task label itself. This trade-off holds regardless of the technique used to learn the feature mappings that produce these representations. We empirically validate this result for a wide range of learning techniques, model architectures, and datasets.
**footnotetext: Equal contribution
@@ 4. IV)
Related Work. Prior works refer to the LPP but so far only study the problem of attribute obfuscation for a single, fixed sensitive attribute. For instance, Song & Shmatikov (2019) experimentally observe that current censoring techniques can only prevent a model from learning a sensitive attribute at the cost of utility loss. Their claim that “overlearning [of sensitive attributes] is intrinsic” (Song & Shmatikov, 2019) to machine learning models is derived from empirical observations on a single learning task and sensitive attribute for a small set of censoring techniques. Zhao et al. (2020) formally derive a lower bound on the trade-off between a model’s performance on its intended learning task and hiding a fixed sensitive attribute.
While these results characterise the trade-off between a representation’s utility and leakage with respect to a specific attribute, they do not actually evaluate whether it is possible to learn useful representations that fulfil the LPP. They focus on a single attribute, while the LPP demands that the representations shared with the service provider should prevent inference of any information other than the intended task (Melis et al., 2019; Brown et al., 2022). Furthermore, neither Song & Shmatikov (2019) nor Zhao et al. (2020) consider that useful representations must at least leak any sensitive information already revealed by the task label itself. As we argue in Section 3, such fundamental leakage is crucial to consider when formalising LPL.
@@ 2 Problem Setup
Assumptions on the Data Distribution. To make our formal analyses tractable, we assume that the spaces $\mathbb{X},\mathbb{Y},\mathbb{S}$ are discrete and finite; and that the data domain is non-trivial ($|\mathbb{X}|>1$ and $|\mathbb{Y}|>1$) and has full support. We also assume that the space of inputs is a subset of the space of sensitive attributes, $\mathbb{X}\subset\mathbb{S}$. A sensitive “attribute” $S$ could be the input $X$ in its entirety. We further make the following assumption about the data distribution:
@@
To protect against such data misuse, users share a representation of their data that restricts information leakage to the necessary minimum for the intended purpose (see Fig. 1). For instance, in the previous example, a representation that only reveals features relevant to recognising a user’s identity. The goal hence is to find a feature mapping $f_{E}(X)=Z$ that maps input $X\in\mathbb{X}$ to a representation $Z\in\mathbb{Z}$ that has high utility for the intended task $Y$ but prevents inference of any other data attribute $S$. In the next section, we formalise this goal and characterise its feasibility.
@@ 3.1 Utility Measure
The representation shared with the service provider should be useful for the intended purpose, i.e., contain features relevant to correctly infer task label $Y$. One common way to formalise this objective is the mutual information between the representations and the task label $I(Y;Z)=I(Y;f_{E}(X))$ (Alemi et al., 2016). The higher the information between $Z$ and $Y$, the more useful the representation will be for the intended task. We make this notion more general bu using a variant of mutual information known as Arimoto’s $\alpha$-information (Arimoto, 1977), which we denote as $I_{\alpha}(Y;Z)$. In particular, we consider two relevant instantiations of $\alpha$: $\alpha=1$ and $\alpha=\infty$. In the case that $\alpha=1$, Arimoto’s information is equal to classical Shannon’s mutual information $I_{\alpha}(Y;Z)=I(Y;Z)$, described previously. In the case that $\alpha=\infty$, Arimoto’s information is the multiplicative gain in accuracy of predicting the task label $Y$ from the learned representations over baseline guessing (Liao et al., 2019):
@@ 3.3 Strawman Approach: Unconditional Least-Privilege Principle
Given a data distribution $P_{X,Y}$, a feature map $f_{E}(X)=Z$ satisfies the unconditional LPP with parameter $\gamma$ if for any attribute $S\neq Y$ which follows the Markov chain SXZS-X-Z, the attribute inference gain is bounded:
@@ 2. II)
We provide a full proof in Appendix A. This result implies that, whenever a representation has a certain utility for the learning task with $I_{\alpha}(Y,Z)>\gamma$, there exists a sensitive attribute for which an adversary’s inference gain is at least as large with $I_{\infty}(S,Z)\geq\gamma$. In fact, under Definition 1, it is easy to construct this attribute to be infinitesimally close to the task label $Y$ but not quite match it. In the next section, we provide an alternative, less literal, formalisation of the LPP that captures the requirement $S\neq Y$ yet precludes these cases.
@@ 3.4 Formalisation of the Least-Privilege Principle
Given a data distribution $P_{X,Y}$, a feature map $f_{E}(X)=Z$ satisfies the LPP with parameter $\gamma$ if for any attribute $S$ which follows the Markov chain S(X,Y)ZS-(X,Y)-Z, the attribute inference gain from observing $(Z,Y)$ over the fundamental leakage is bounded:
@@
LPP and Utility Trade-Off. We now study whether we can find a feature map that achieves the LPP in Definition 2, and simultaneously has good utility for the intended learning task. We show the general trade-off between a representation’s utility and the LPP requirement:
@@ 2. II)
One way to see why the trade-off holds is that there exists a set of attributes that reveal maximum possible information about $X$ from $Z$ (see Appendix A for a formal description). Intuitively, we would expect that by allowing for fundamental leakage, i.e., by conditioning on $Y$, we might reduce maximal leakage in those cases where $Y$ is one of these maximally revealing attributes. However, as we show in Appendix A, strict positivity of the posterior distribution (Assumption A) prevents the task label $Y$ to be such a maximally revealing attribute and the least-privilege and utility trade-off thus remains the same as under unconditional LPP (Theorem 1).
An important implication of Theorem 2 is that perfect LPP with $\gamma=0$, where the feature representations do not allow for any inference gain over fundamental leakage, is only possible when the representations are constant or completely random:
Corollary 1.
[…]
@@
The question of finding a feature representation that maximises utility but satisfies the LPP posed by Theorem 2 is a variant of the generalised conditional entropy bottleneck (CEB) problem (Fischer, 2020), which, in turn is a variant of the standard information bottleneck problem (Asoodeh & Calmon, 2020; Tishby et al., 2000). The difference with respect to the original formulation of the CEB problem is that the CEB constrains the conditional mutual information between the input and the representation $I(X;Z\mid Y)$ rather than its maximal leakage $\mathcal{L}(X\rightarrow Z\mid Y)$.
Unlike in the CEB, we require a leakage measure that captures the original claim from prior work that it is possible to learn representations that protect against any harmful inferences, i.e., protect even against worst-case inference adversaries. Maximal leakage satisfies this requirement and in addition has a direct operational meaning in terms inference gain (see Eq. 4).
0112201122Feasible regionAttribute inference gain, $\gamma=\sup_{S}\log\frac{\Pr[S=\hat{S}(Z,Y)]}{\Pr[S=\hat{S}(Y)]}$Task utility, $I_{\alpha}(Y;Z)$
@@
Traditionally, a power of $e$ instead of $2$ is used, but we use the latter for consistency with our previous definitions.
Given this definition, we show that LDP implies LPP (see proof in Appendix A) :
@@
An important implication is that a feature map that satisfies the LPP comes with the same utility trade-off as one that satisfies LDP. Thus, LPL cannot provide an alternative to differential privacy at a lower utility cost as envisioned by prior work (Osia et al., 2018; Melis et al., 2019).
@@ 4 Empirical Evaluation
$\tilde{\Pr}[S=s\mid Y=y]=\frac{\sum_{i=1}^{A}\mathbf{1}[s_{i}=s,y_{i}=y]}{\sum_{i=1}^{A}\mathbf{1}[y_{i}=y]},$
and outputs a guess according to $\smash{\hat{S}(Y=y)=\arg\max_{s}\tilde{\Pr}[S=s\mid Y=y]}.$ The features adversary $\hat{S}(Z,Y)$ is given black-box access to the feature mapping of the trained model. To collect a train set, the features adversary submits each example $x_{i}\in D_{A}$ to the model and receives back its representation $z_{i}=f_{E}(x_{i})$. The adversary then trains a Random Forest classifier with $50$ decision trees on the collected samples $(z_{i},y_{i},s_{i})$ to estimate $\smash{\tilde{\Pr}[S=s\mid Z=z,Y=y]}$. We opt for Random Forest as an attack model based on its superior performance over other classifiers we tested.
@@
Learning Techniques. We implement two learning techniques: (1) standard ERM with SGD, and (2) attribute censoring to learn representations that hide a given sensitive attribute. For censoring, we use the gradient reversal strategy (GRAD) introduced by Raff & Sylvester (2018) with a censoring parameter of $100$. We choose GRAD because it effectively hides the chosen sensitive attribute without large drops in model performance (Zhao et al., 2020), and, unlike other censoring techniques, can be applied to any model architecture. In Appendix B, we show equivalent results for another learning technique that aims to hide sensitive information, adversarial representation learning.
@@ 4.1 The Potential Harms of Fundamental Leakage
Figure 3: Fundamental leakage: the task label reveals information about other data attributes; which might not be obvious to data subjects. Attribute inference gain of the label-only adversary (left) and pairwise Pearson’s correlation between attributes (right). In the LFWA+ dataset, the ‘Attractive’ label is highly correlated with the perceived gender. Thus, predicting the ‘Attractive’ label will reveal information about gender.
Fig. 3 (left) shows the adversary’s gain $\smash{\smash{\tilde{I}_{\infty}}(S;Y)}$ in predicting sensitive attribute $S$ from $Y$ for pair-wise combinations of $12$ learning tasks and sensitive attributes. This gain represents the fundamental leakage of a system that fulfils its intended purpose, i.e., that produces accurate task labels for the chosen learning task. The matrix in Fig. 3 (right) shows the absolute pairwise Pearson’s correlation coefficient between attribute labels. Fig. 3 shows that, as expected, a strong linear relationship between the learning task and the sensitive attribute targeted by the adversary leads to a large fundamental leakage. For instance, ‘Attractive’ and ‘Male’ are negatively correlated with a correlation coefficient of 0.3094-0.3094. This is already enough to increase the adversary’s gain in inferring sensitive attribute ‘Male’ when the learning task is ‘Attractive’. Although the increase is small in this case, it illustrates how inferences due to a task’s fundamental leakage can not only be counterintuitive, but also reveal information that could lead to harm (in this case, discrimination due to gender).
@@ 4.2 The Least-Privilege And Utility Trade-Off
Figure 4: If the model-generated representations have utility for the task (right), there exists a sensitive attribute with an even higher inference gain for the adversary (left, red means more leakage). This holds for both standard ERM (top) and attribute censoring (bottom) where we censor the attribute with highest leakage in the respective ERM model (marked as ). Censoring has a ‘whack-a-mole’ effect: as we censor one attribute, leakage of another attribute increases.
@@
Across Learning Tasks. In Fig. 4 (top), we see that in the LFWA+ dataset the features learned by a model that predicts attribute ‘Smiling’ increase the adversary’s inference gain for attribute ‘White’. Different tasks result in a high leakage for different attributes, e.g., ‘OvalFace’ reveals a lot of additional information about gender, while ‘HeavyMakeup’ is indicative of ‘PaleSkin’ and ‘Attractive’ but does not have much influence on the inference power of other attributes. Importantly, for every task, there is at least one sensitive attribute with $\Delta_{\mathsf{Adv}}>0$. These results confirm that the features learned by a model trained to perform well on its learning task reveal more information than a task’s fundamental leakage and thus violate the LPP.
@@
Across Datasets and Model Architectures. Besides images, tabular data is another type of data for which representation sharing is commonly suggested as a technique to limit unintended information leakage (Zhao et al., 2020; Arik & Pfister, 2021; Creager et al., 2019; Raff & Sylvester, 2018). In Figs. 5 and 9 in Appendix B, we show that the trade-off also holds for this type of data (and the associated model architectures) using two tabular datasets: it is not possible to find feature representations that have high utility for a prediction task but fulfil the LPP.

### v2

@@ Abstract
The promise of least-privilege learning – to find feature representations that are useful for a learning task but prevent inference of any sensitive information unrelated to this task – is highly appealing. However, so far this concept has only been stated informally. It thus remains an open question whether and how we can achieve this goal. In this work, we provide the first formalisation of the least-privilege principle for machine learning and characterise its feasibility. We prove that there is a fundamental trade-off between a representation’s utility for a given task and its leakage beyond the intended task: it is not possible to learn representations that have high utility for the intended task but, at the same time prevent inference of any attribute other than the task label itself. This trade-off holds under realistic assumptions on the data distribution and regardless of the technique used to learn the feature mappings that produce these representations. We empirically validate this result for a wide range of learning techniques, model architectures, and datasets.
Privacy, Trustworthy Machine Learning
@@ 4. IV)
Related Work. Prior works refer to the LPP but so far only study the problem of attribute obfuscation for a single, fixed sensitive attribute. For instance, Song & Shmatikov (2019) experimentally observe that current censoring techniques can only prevent a model from learning a sensitive attribute at a cost in model utility. Their claim that “overlearning [of sensitive attributes] is intrinsic” (Song & Shmatikov, 2019) to machine learning models is derived from empirical observations on a single learning task and sensitive attribute for a small set of censoring techniques. Zhao et al. (2020) formally derive a lower bound on the trade-off between a model’s performance on its intended learning task and hiding a fixed sensitive attribute.
While these results characterise the trade-off between a representation’s utility and leakage with respect to a specific attribute, they do not actually evaluate whether it is possible to learn useful representations that fulfil the LPP. They focus on a single attribute, whereas the LPP demands that the representations shared with the service provider should prevent inference of any information other than the intended task (Melis et al., 2019; Brown et al., 2022). Furthermore, neither Song & Shmatikov (2019) nor Zhao et al. (2020) consider that useful representations must at least leak any sensitive information already revealed by the task label itself. As we argue in Section 3, such fundamental leakage is crucial to consider when formalising LPL.
@@ 2 Problem Setup
Assumptions on the Data Distribution. To make our formal analyses tractable, we assume that the spaces $\mathbb{X},\mathbb{Y},\mathbb{S}$ are discrete and finite; and that the data domain is non-trivial ($|\mathbb{X}|>1$ and $|\mathbb{Y}|>1$) and has full support. We further make the following assumption about the data distribution:
@@
To prevent such data misuse, users share a representation of their data that restricts information leakage to the necessary minimum for the intended purpose (see Fig. 1). For instance, in the previous example, a representation that only reveals features relevant to recognising a user’s identity. The goal hence is to find a possibly randomized feature mapping $f_{E}(\cdot)$ that maps input $X$ to a representation $Z=f_{E}(X)$ that both has high utility for the intended task $Y$ but prevents inference of any other data attribute $S$. We formalise this goal and characterise its feasibility in the next section.
@@ 3.1 Utility Measure
The representation shared with the service provider should be useful for the intended purpose, i.e., contain features relevant to correctly infer task label $Y$. One common way to formalise this objective is the mutual information between the representations and the task label $I(Y;Z)=I(Y;f_{E}(X))$ (Alemi et al., 2016). The higher the information between $Z$ and $Y$, the more useful the representation will be for the intended task. We make this notion more general by using a variant of mutual information known as Arimoto’s $\alpha$-information (Arimoto, 1977), which we denote as $I_{\alpha}(Y;Z)$. In particular, we consider two relevant instantiations of $\alpha$: $\alpha=1$ and $\alpha=\infty$. In the case that $\alpha=1$, Arimoto’s information is equal to classical Shannon’s mutual information $I_{\alpha}(Y;Z)=I(Y;Z)$, described previously. In the case that $\alpha=\infty$, Arimoto’s information is the multiplicative gain in accuracy of predicting the task label $Y$ from the learned representations over baseline guessing (Liao et al., 2019):
@@ 3.3 Strawman Approach: Unconditional Least-Privilege Principle
Given a data distribution $P_{X,Y}$, a feature map $f_{E}(X)=Z$ satisfies the unconditional LPP with parameter $\gamma$ if for any attribute $S\neq Y$ which follows the Markov chain $S-X-Z$, the attribute inference gain is bounded:
@@ 2. II)
We provide a proof of this and all of the following formal statements in Appendix A. This result implies that whenever a representation has a certain utility for the learning task with $I_{\alpha}(Y,Z)>\gamma$, there exists a sensitive attribute for which an adversary’s inference gain is at least as large with $I_{\infty}(S,Z)\geq\gamma$. In fact, under Definition 1, it is easy to construct this attribute to be infinitesimally close to the task label $Y$ but not quite match it. In the next section, we provide an alternative formalisation of the LPP that captures the requirement $S\neq Y$ yet precludes these cases.
@@ 3.4 Formalisation of the Least-Privilege Principle
Given a data distribution $P_{X,Y}$, a feature map $f_{E}(X)=Z$ satisfies the LPP with parameter $\gamma$ if for any attribute $S$ which follows the Markov chain $S-(X,Y)-Z$, the attribute inference gain from observing $(Z,Y)$ over the fundamental leakage is bounded:
@@
Perfect LPP. We first analyse the feasibility of perfect LPP with $\gamma=0$. Perfect LPP implies that the representation $Z$ leaks strictly no more information about the data $X$ than already revealed by the task label $Y$ itself. As we show next, perfect LPP is possible only under a restrictive assumption on the data distribution.
Proposition 1.
There exists a feature map $f_{E}(X)$ that fulfils the LPP with $\gamma=0$ if and only if we have a Markov chain $X-Y-Z$.
One class of data distributions that fulfils this condition and where perfect LPP is feasible is any distribution in which the intended task label $Y$ is a deterministic function of the input data $X$ (see, e.g., Fischer (2020)):
Corollary 1.
If $Y=g(X)$ for some function $g:\mathbb{X}\rightarrow\mathbb{Y}$, there exists $f_{E}(X)$ which achieves perfect LPP with $\gamma=0$.
As we discuss in Section 2, in many realistic scenarios, however, there is inherent uncertainty about an example’s true label due to, for instance, label noise. Under Assumption A, perfect LPP is unattainable by non-trivial feature maps:
Corollary 2.
[…]
LPP and Utility Trade-Off. We now study whether we can find a feature map that achieves the LPP in Definition 2 with $\gamma>0$, and simultaneously has good utility for the intended learning task. We show the general trade-off between a representation’s utility and the LPP requirement:
@@ 2. II)
One way to see why the trade-off holds is through the lens of attributes that reveal maximum possible information about $X$ from $Z$ (see Appendix A for a formal description). Intuitively, we would expect that by allowing for fundamental leakage, i.e., by conditioning on $Y$, we might reduce maximal leakage in those cases where $Y$ is a maximally revealing attribute. However, as we show in Appendix A, strict positivity of the posterior distribution (Assumption A) prevents the task label $Y$ from being a maximally revealing attribute, thus the least-privilege and utility trade-off remains the same as under unconditional LPP (Theorem 1).
@@
Our formalisation of the LPP is closely related to the Minimum Necessary Information (MNI) principle introduced by Fischer (2020). The MNI criterion demands that a representation $Z$ should maximise the mutual information $I(Y;Z)$ while minimising the conditional information $I(X;Z\mid Y)$. The LPP and MNI differ in their leakage measure, but are equal in the optimal point of $\gamma=0$ in which $I(X;Z\mid Y)=\mathcal{L}(X\rightarrow Z\mid Y)=0$.
The question of finding a feature representation that maximises utility but satisfies the LPP posed by Theorem 2 can be seen as a variant of the conditional entropy bottleneck (CEB) problem (Fischer, 2020), which, in turn is a variant of the standard information bottleneck problem (Asoodeh & Calmon, 2020; Tishby et al., 2000). The CEB problem aims to recover a representation that satisfies the MNI principle.
Unlike in the MNI and CEB, we require a leakage measure that captures the original claim from prior work that it is possible to learn representations that protect against any harmful inferences, i.e., protect even against worst-case inference adversaries. Maximal leakage satisfies this requirement and in addition has a direct operational meaning in terms of inference gain (see Eq. 4).
Other works, such as Mireshghallah et al. (2021) or Maeng et al. (2024), differ not only in their leakage measure but crucially do not consider a task’s fundamental leakage, i.e., do not define the trade-off in terms of the conditional leakage $I_{\alpha}(S,Z\mid Y)$. In addition, Mireshghallah et al. (2021) and other formalisations of related information-theoretic problems, such as the Privacy Funnel (Makhdoumi et al., 2014; Asoodeh & Calmon, 2020), only consider a fixed sensitive attribute which does not adequately capture the least-privilege requirement.
00111122220011112222Feasible regionAttribute inference gain, $\gamma=\sup_{S}\log\frac{\Pr[S=\hat{S}(Z,Y)]}{\Pr[S=\hat{S}(Y)]}$ask utility, $I_{\alpha}(Y;Z)$
@@
Conventionally, a power of $e$ instead of $2$ is used, but we use the latter for consistency with our previous definitions.
Given this definition, we show that LDP implies LPP:
@@
An important implication is that a feature map that satisfies the LPP comes with the same utility trade-off as one that satisfies LDP. Thus, in the setting of Theorem 2, LPL cannot provide an alternative to differential privacy at a lower utility cost as envisioned by prior work (Osia et al., 2018; Melis et al., 2019).
@@ 4 Empirical Evaluation
$\tilde{\Pr}[S=s\mid Y=y]~{}\triangleq~{}\frac{\sum_{i=1}^{A}\mathbf{1}[s_{i}=s,y_{i}=y]}{\sum_{i=1}^{A}\mathbf{1}[y_{i}=y]},$
and outputs a guess according to:
$\smash{\hat{S}(Y=y)=\arg\max_{s}\tilde{\Pr}[S=s\mid Y=y]}.$
The features adversary $\hat{S}(Z,Y)$ is given black-box access to the feature mapping of the trained model. To collect a train set, the features adversary submits each example $x_{i}\in D_{A}$ to the model and receives back its representation $z_{i}=f_{E}(x_{i})$. The adversary then trains a Random Forest classifier with $50$ decision trees on the collected samples $(z_{i},y_{i},s_{i})$ to estimate $\smash{\tilde{\Pr}[S=s\mid Z=z,Y=y]}$. We opt for Random Forest as an attack model based on its superior performance over other classifiers we tested.
@@
Learning Techniques. We implement two learning techniques: (1) standard ERM with SGD, and (2) attribute censoring to learn representations that hide a given sensitive attribute. In our main experiment, we use the gradient reversal strategy (GRAD) introduced by Raff & Sylvester (2018) for censoring with a censoring parameter of $100$. We choose GRAD because it effectively hides the chosen sensitive attribute without large drops in model performance (Zhao et al., 2020), and, unlike other censoring techniques, can be applied to any model architecture. In Appendix B, we show equivalent results for another learning technique that aims to hide sensitive information, adversarial representation learning.
@@ 4.1 The Potential Harms of Fundamental Leakage
Figure 3: Fundamental leakage: the task label reveals information about other data attributes, which might not be obvious to data subjects. Attribute inference gain of the label-only adversary (left) and pairwise Pearson’s correlation between attributes (right). In the LFWA+ dataset, the ‘Attractive’ label is highly correlated with the perceived gender. Thus, predicting the ‘Attractive’ label will reveal information about gender.
Fig. 3 (left) shows the adversary’s gain $\smash{\smash{\tilde{I}_{\infty}}(S;Y)}$ in predicting sensitive attribute $S$ from $Y$ for pair-wise combinations of $12$ learning tasks and sensitive attributes. This gain represents the fundamental leakage of a system that fulfils its intended purpose, i.e., that produces accurate task labels for the chosen learning task. The matrix in Fig. 3 (right) shows the absolute pairwise Pearson’s correlation coefficient between attribute labels. Fig. 3 shows that, as expected, a strong linear relationship between the learning task and the sensitive attribute targeted by the adversary leads to a large fundamental leakage. For instance, ‘Attractive’ and ‘Male’ are negatively correlated with a correlation coefficient of $-0.3094$. This is already enough to increase the adversary’s gain in inferring sensitive attribute ‘Male’ when the learning task is ‘Attractive’. Although the increase is small in this case, it illustrates how inferences due to a task’s fundamental leakage can not only be counterintuitive, but also reveal information that could lead to harm (in this case, discrimination due to gender).
@@ 4.2 The Least-Privilege And Utility Trade-Off
[…]
@@
Across Learning Tasks. In Fig. 4 (top), we see that in the LFWA+ dataset the features learned by a model that predicts attribute ‘Smiling’ increase the adversary’s inference gain for attribute ‘White’. Different tasks result in a high leakage for different attributes, e.g., ‘OvalFace’ reveals a lot of additional information about gender, while ‘HeavyMakeup’ is indicative of ‘PaleSkin’ and ‘Attractive’ but does not have much influence on the inference power of other attributes. Importantly, for every task, there is at least one sensitive attribute with $\Delta_{\mathsf{Adv}}>0$. These results confirm that the features learned by a model trained to perform well on its learning task reveal more information than a task’s fundamental leakage and thus violate the LPP. Fig. 5 in Appendix B shows these results in terms of the adversary’s absolute inference gain $\smash{\tilde{I}_{\infty}}(S;Z\mid Y)$. It confirms that the adversary’s absolute inference gain increases with a model’s performance on its intended task.
@@
Across Datasets and Model Architectures. Besides images, tabular data is another type of data for which representation sharing is commonly suggested as a technique to limit unintended information leakage (Zhao et al., 2020; Arik & Pfister, 2021; Creager et al., 2019; Raff & Sylvester, 2018). In Figs. 6 and 10 in Appendix B, we show that the trade-off also holds for this type of data (and the associated model architectures) using two tabular datasets: it is not possible to find feature representations that have high utility for a prediction task but fulfil the LPP.
@@
5 Limitations and Future Work
Our results hold under a mild assumption on the data distribution (Assumption A) and for a specific operational measure of leakage (Section 3.2). By relaxing these two aspects, it might be possible to achieve a more favourable trade-off than what we find.
First, the trade-off could possibly be less stringent for specific data distributions for which Assumption A does not hold. In such cases, it might be possible that for some learning tasks, there exists a representation that provides some utility gain and at the same time constrains the maximal leakage, i.e., that simultaneously fulfils the two conditions stated in Theorem 2. It could be a direction for future work to formally analyse whether such representations are achievable for distributions that do not satisfy Assumption A.
Second, the leakage measure used to formalise the LPP could be relaxed by restricting the set of possible inference targets to a less rich class of distributions. For instance, the class of attributes targeted by the adversary could be limited to those “efficiently identifiable” (Hébert-Johnson et al., 2018). However, such modifications would require a different analytical toolbox and deviate from the promise of LPL to leak “nothing else” (Brown et al., 2022) but the intended task. In addition, although the leakage measure considered here might seem strong, it is already significantly weaker than standard privacy notions, such as LDP, as it measures attribute inference success in expectation over the population as opposed to worst-case risks.